# Temporal Feature Matters: A Framework for Diffusion Model Quantization

### v2

@@ 1 Introduction
To tackle temporal feature disturbance, we first find that the modules generating temporal features are independent of the sampling data and thus treat the whole modules as the Temporal Information Block (TIB). However, existing methods [18, 42, 72, 77, 74] do not separately optimize this block during the quantization process, causing temporal features to overfit to limited calibration data. Additionally, since the maximum time-step for denoising is a finite positive integer, the temporal feature and the activations during its generation form a finite set. Therefore, the optimal approach is to optimize each element in this set individually. Inspired by these observations and analyses, we propose two temporal feature maintenance techniques: 1) TIB-based Maintenance: Based on TIB, a novel quantization reconstruction approach, Temporal Information-aware Reconstruction (TIAR), is devised. It aims to reduce temporal feature error as the optimization objective while isolating the network’s components related to sampled data during weight adjustment. Besides, we also introduce a calibration strategy, Finite Set Calibration (FSC), for the finite set of the temporal feature and activation during its generation. 2) Cache-based Maintenance: Further leveraging the independent and finite nature of the temporal feature, we directly pre-compute these features, and then optimize and cache their quantized versions to reduce the disturbance. In detail, this novel design only requires reloading cached features, avoiding online generation during inference, which enhances latency in specific cases. Moreover, we employ our Disturbance-aware Selection taking advantage of both maintenance approaches to further eliminate temporal feature error. In addition, only one offline generation process for fine-grained maintenance strategies chosen ensured the efficiency of our selection scenario. Finally, evaluating on multiple datasets, diverse tasks, advanced models, and different hardware, our novel framework towards reducing temporal feature disturbance achieved nearly lossless model compression with high inference speed.
@@
This paper extends our conference version [23] in several aspects. 1) We compare and fine-grainedly analyze the sensitivity and quantization-induced disturbance between temporal features and other non-temporal features. 2) We propose Cache-based Maintenance parallel to TIB-based Maintenance with high efficiency. It can achieve comparable performance and even lower inference costs in some cases. 3) We propse Disturbance-aware Selection fine-grainedly harness both maintenance strategies to further reduce disturbance in temporal features. 4) We apply our framework to advanced Text-to-Image (T2I) models, e.g., SD-XL [64] and SD-XL-turbo [71]. 5) We deploy our quantized model employing CUTLASS [32] with our customized kernel on Nvidia H800 GPU and OpenVino [25] on Intel® Xeon® Gold 6448Y Processor to show significant acceleration. 6) We provide comprehensive experiments and more ablative studies to investigate the effectiveness of our methods.
@@ 4.1 Disturbance Observations
Figure 3: (Left) Temporal feature disturbance. The green inflection points highlight the significant phenomenon of temporal feature disturbance. (Right) Temporal information mismatch ($i=11$). The coordinates of the inflection points on the blue curve can denoted as $(t,t+\delta_{t,i})$. It indicates $\widehat{\mathbf{X}_{t,i}^{temporal}}$ exhibits the highest similarity with $\mathbf{X}_{t+\delta_{t,i},i}^{temporal}$.
As illustrated in Fig. 3 (Left), quantization induces notable errors in temporal features, far exceeding those in non-temporal ones. We refer to this phenomenon, characterized by substantial temporal feature errors within diffusion models, as temporal feature disturbance.
@@ 4.3 Inducement Analyses
TABLE I: FID, sFID and SQNR on LSUN-Bedrooms $256\times 256$ [85] for LDM-4. Prev represents BRECQ. Freeze denotes our trial.
@@ 4.4 Quantization Framework
Figure 5: Overview of the proposed framework. (a) TIB-based Maintenance: Based on a Temporal Information Block, we enable Temporal Information-aware Reconstruction and Finite Set Calibration. (b) Cache-based Maintenance: We directly quantize these temporal features and reload them for inference. (c) Disturbance-aware Selection: We select the appropriate maintenance through temporal feature error for further improvement. This framework achieves the maintenance of temporal features and yields promising results.
@@ 4.4.1 TIB-based Maintenance
Temporal Information Block (TIB). Based on the inducements, it is crucial to meticulously separate the reconstruction and calibration process for each embedding layer and Residual Bottleneck Block to enhance quantized model performance. Considering the unique structure of the UNet, we consolidate all embedding layers and time embed into a unified Temporal Information Block (TIB), which can be denoted as $\{g_{i}(h(\cdot))\}_{i=0,\ldots,n}$ (see Fig. 5 (a)).
Temporal Information-aware Reconstruction. Based on the Temporal Information Block (TIB), we introduce the Temporal Information-aware Reconstruction (TIAR) to address the first inducement. The optimization goal for the block during the reconstruction phase is captured by the following loss function
$\mathcal{L}_{TIAR}=\sum_{i=0}^{n}\|g_{i}(h(t))-\widehat{g}_{i}(\widehat{h}(t))\|_{F}^{2},$ (11)
where $\widehat{h}(\cdot)$ and $\widehat{g}_{i}(\cdot)$ represent the quantized counterparts of $h(\cdot)$ and $g_{i}(\cdot)$, respectively. This reconstruction strategy focuses on adjusting weights to pursue a minimal disturbance for temporal features.
Finite Set Calibration. To address the challenge posed by the wide span of activations within a finite set for the second inducement, we propose Finite Set Calibration (FSC) for activation quantization. This strategy employs $T$ sets of quantization parameters for every activation within all the embedding layers and time embed, such as $\{(s_{T},z_{T}),\ldots,(s_{1},z_{1})\}$ for activation $\mathbf{X}$. At time-step $t$, the quantization function for the $\mathbf{X}$ can be expressed as:
@@
where $s_{t}$ and $z_{t}$ are the corresponding scale and zero offset. To be noted, the calibration target for these activations is also aligned with the output of the TIB. Besides that, we find that Min-max [58] can achieve satisfactory results with high efficiency (more evidence in Sec. 5.4.2) for range estimation. In Sec. 5.3.1, we perform a detailed analysis to demonstrate that our method incurs a negligible extra cost for inference.
@@ 4.4.2 Cache-based Maintenance
Further considering sample independence and finitude of the temporal feature, the feature associated with each $t$ and $i$ remains constant and can therefore be pre-computed offline. This allows us to directly optimize the quantization parameters for these pre-computed features and cache the quantized counterparts with the parameters to address the issues. The objective for all $t$ and $i$ can be formulated as follows:
@@
where $s_{t,i}$ and $z_{t,i}$ denote quantization scale and zero offset for the pre-gained $\mathbf{X}_{t,i}^{temporal}$. Moreover, we employ LSQ [11] to optimize every objective separately, which can be done in just a few minutes. During inference, this strategy enables us to get the temporal feature by reloading the cached items. In addition, we find that this approach can help improve latency. More detailed analysis can be found in Sec. 5.3.2.
@@ 4.4.3 Disturbance-aware Selection
We visualize the comparison of the temporal feature disturbance between the two maintenance approaches in Fig. 6, where the grid value of $(t,i)$ represents the temporal feature error [^6] Here we employ $\mathcal{L}_{t,i}^{temporal}$ in Eq. (13), instead of that in Eq. (8) as temporal feature error, since over $82\%$ of cosine similarity with our two maintenance approaches are almost exactly equal to 1, which shows the remarkable effect of our methods. as follows:
@@
Given the discrepancy in temporal feature errors across different $t$ and $i$ in our two methods, we propose Disturbance-aware Selection to better mitigate the temporal feature disturbance. Specifically, when $\tau_{t,i}<1$, we employ TIB-based Maintenance to obtain $\widehat{\mathbf{X}_{t,i}^{temporal}}$. Conversely, we opt for Cache-based Maintenance. For efficiency, our selection procedure can be done offline with only one image generation pass. To validate its performance, Fig. 7 demonstrates that our straightforward but effective selection strategy successfully leverages the strengths of both maintenance approaches.
@@ 5 Experiments
We organize the experiments as follows: In Sec. 5.1, we first demonstrate implementation details. Then, we exhibit the outstanding performance gained by our framework in Sec. 5.2. Further, we also deploy our quantized models on various hardware to investigate their inference efficiency in Sec. 5.3. Last, we conduct comprehensive ablative studies for the proposed framework in Sec. 5.4.
@@ 5.2 Performance Comparison
In the experiments conducted on the LDM with DDIM sampler [75], we maintain the same experimental settings as presented in  [67], including the number of steps, variance schedule, and classifier-free guidance scale (denoted by eta and cfg in the following, respectively). As shown in Tab. II, the FID performance differences relative to the full precision (FP) model are all within 0.7 for all settings employing DS. Specifically, on the CelebA-HQ $256\times 256$ dataset, DS exhibits an FID reduction of 6.74, an sFID reduction of 6.79, and an SQNR increasing of 4.09 in the w4a8 setting compared to the current algorithms. It is noticeable that existing methods, whether in 4-bit or 8-bit, show significant performance degradation when compared to the FP model on face datasets like CelebA-HQ $256\times 256$ and FFHQ $256\times 256$, whereas our TM/CM/DS shows almost no performance degradation. Importantly, DS also achieves significant performance improvement on the LSUN-Bedrooms $256\times 256$ in the w4a8 setting, with FID and sFID reductions of 2.33 and 7.67 compared to PTQD [18], respectively. Regarding LDM-8 on LSUN-Churches $256\times 256$, we attribute the moderate improvement, compared to other datasets. We believe that the use of the LDM-8 model with a downsampling factor of 8 may be more quantization-friendly. Existing methods have already achieved satisfactory results on this dataset. Nonetheless, our method still approaches the performance of the FP model more closely than existing methods.
@@
Unconditional generation.
TABLE IV: Quantization results for text-guided image generation with Stable Diffusion-v1-4, SD-XL-base-1.0, and SD-XL-turbo on MS-COCO captions.
[…]
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
TABLE V: Quantization results for unconditional image generation with class-conditional image generation with LDM-4 on ImageNet $256\times 256$.
@@
Besides the experiments on LDM, We have also conducted experiments with DDPM on CIFAR-10 $32\times 32$. As shown in Tab. III, our methods still achieve comprehensive improvements in terms of IS, FID, and SQNR compared to the existing advanced mthods. However, due to the lower resolution and simplicity of the images in this dataset, existing methods show minimal performance degradation, so the results we obtain may not be as pronounced.
Class-conditional generation. On the ImageNet $256\times 256$ dataset, we employed a denoising process with 20 steps DDIM sampler [75], setting eta and cfg to 0.0 and 3.0, respectively. In Tab. V, compared to PTQD, our method achieves a FID reduction of 2.09 on w8a8, and a 5.35 sFID decrease on w4a8. Under all the conditions with DS, we observe an improvement of over 9 in IS. Particularly noteworthy is that, across various quantization settings, our method consistently achieved lower FID and higher SQNR compared to the FP model.
Text-guided generation.
In this experiment, we generate high-resolution images of $512\times 512$ pixels using Stable Diffusion-v1-4 with 50 steps PLMS sampler [48], and SD-XL-turbo with only a single step Euler sampler [30]. Moreover, higher resolution images of $1024\times 1024$ are also generated from SD-XL-base-1.0 with 50 steps Euler sampler. cfg is fixed to the default 7.5 for Stable Diffusion and SD-XL as the trade-off between generation quality and diversity. However, we do not use classifier-free guidance [22] to save memory, the same as the original paper [71]. In Tab. IV, compared to Q-Diffusion, our approach achieves an FID reduction of 1.20, 5.61, and 5.50 under the w4a8 configuration for Stable Diffusion, SD-XL, and SD-XL-turbo, respectively. Notably, our FID on w8a8 and sFID on w4a32 are even lower than those of the full precision model. The table also shows our significantly lower quantization error than Q-Diffusion. For example, DS obtains an 8.16 SQNR upswing for w4a8 Stable Diffusion. Benefiting from this, the visual effect of our generated images  [^8] More qualitative comparisons can be found in the supplementary materials. is closer to the FP model with better quality than Q-Diffusion (examples in Fig. 8). All of these clues strongly reveal the value of maintaining temporal features.

### v3

@@ 1 Introduction
To tackle temporal feature disturbance, we first find that the modules generating temporal features are independent of the sampling data and thus treat the whole modules as the Temporal Information Block (TIB). However, existing methods [19, 44, 79, 85, 81] do not separately optimize this block during the quantization process, causing temporal features to overfit to limited calibration data. Additionally, since the maximum time-step for denoising is a finite positive integer, the temporal feature and the activations during its generation form a finite set. Therefore, the optimal approach is to optimize each element in this set individually. Inspired by these observations and analyses, we propose two temporal feature maintenance techniques: 1) TIB-based Maintenance: Based on TIB, a novel quantization reconstruction approach, Temporal Information-aware Reconstruction (TIAR), is devised. It aims to reduce temporal feature error as the optimization objective while isolating the network’s components related to sampled data during weight adjustment. Besides, we also introduce a calibration strategy, Finite Set Calibration (FSC), for the finite set of the temporal feature and activation during its generation. 2) Cache-based Maintenance: Further leveraging the independent and finite nature of the temporal feature, we directly pre-compute these features, and then optimize and cache their quantized versions to reduce the disturbance. In detail, this novel design only requires reloading cached features, avoiding online generation during inference, which enhances latency in specific cases. Moreover, we employ our Disturbance-aware Selection taking advantage of both maintenance approaches to further eliminate temporal feature error. In addition, only one generation process before deployment can determine the selection results for every single temporal feature, respectively. This highly ensures the efficiency of our selection scenario. Finally, evaluating on multiple datasets, diverse tasks, advanced models, and different hardware, our novel framework for reducing temporal feature disturbance achieved nearly lossless model compression with high inference speed.
@@
This paper extends our conference version [24] in several aspects. 1) We compare and fine-grainedly analyze the sensitivity and quantization-induced disturbance between temporal features and other non-temporal features. 2) We propose Cache-based Maintenance parallel to TIB-based Maintenance with high efficiency. It can achieve comparable performance and even lower inference costs in some cases. 3) We propose Disturbance-aware Selection fine-grainedly harness both maintenance strategies to further reduce disturbance in temporal features. 4) We apply our framework to advanced Text-to-Image (T2I) models, e.g., SD-XL [71], SD-XL-turbo [78], and FLUX.1-Schnell [42]. Moreover, we also apply our methods to the video generation model–OpenSora [103]. 5) We deploy our quantized model employing CUTLASS [33] with our customized kernel on Nvidia H800 GPU and OpenVino [26] on Intel® Xeon® Gold 6448Y Processor to show significant acceleration. Additionally, we also test the efficiency of the quantized model on Nvidia Jetson Nano Orin and iPhone 15 Pro Max. 6) We provide comprehensive experiments and more ablative studies to investigate the effectiveness of our methods.
@@ 2.1 Efficient Diffusion Models
Besides the above-mentioned studies, there are also some works, e.g., SnapFusion [46] and BitsFusion [84] aim to extremely compress the diffusion model with a mix of multiple techniques for deployment on edge devices. However, these methods also require extensive resources for training or fine-tuning.
@@ 4.1 Disturbance Observations
Figure 3: (Left) Temporal feature disturbance. The green inflection points highlight the significant phenomenon of temporal feature disturbance. (Right) Temporal information mismatch ($i=11$). The coordinates of the inflection points on the blue curve can denoted as $(t,t+\delta_{t,i})$. It indicates $\widehat{\mathbf{X}_{t,i}^{temporal}}$ exhibits the highest similarity with $\mathbf{X}_{t+\delta_{t,i},i}^{temporal}$. To be noted, for any point on the polylines in the above figures, assuming its x-axis is $t_{0}\in[1,\ldots,T]$, we pursue $T-t_{0}+1$ denoising processes to obtain its y-axis from the same random noise. Specifically, the model is set to full precision from $T$ to $t_{0}+1$ and only quantized precision at $t_{0}$.
As illustrated in Fig. 3 (Left). quantization induces notable errors in temporal features, far exceeding those in non-temporal ones. We refer to this phenomenon, characterized by substantial temporal feature errors within diffusion models, as temporal feature disturbance.
@@ 4.2 Disturbance Impacts
We define temporal feature disturbance as a significant temporal feature error (Sec. 4.1). Then we show that temporal feature disturbance (i.e., significant temporal feature error) accounts for temporal information mismatch (Sec. 4.2), which leads to trajectory deviation (Sec. 4.2). In Fig. 4, the deviated trajectory can result in noisy images gradually shifting from a normal denoising trajectory. Thus, the image quality degrades severely. Moreover, as reflected in Fig. 2, with temporal feature error increasing [^4] Obviously, temporal feature error which can also be defined by $\mathcal{L}^{temporal}_{t,i}$ in Eq. (13) increases with the increase of noise level $\lambda$., generation quality measured by FID score deteriorates sharper and sharper. Sec. A for a more detailed relationship between temporal feature errors and final generation quality.
@@
Figure 5: Overview of the proposed framework. (a) TIB-based Maintenance: Based on a Temporal Information Block, we enable Temporal Information-aware Reconstruction and Finite Set Calibration for weights and activations quantization, respectively. (b) Cache-based Maintenance: We directly quantize pre-fetched temporal features and cache their quantized version. During inference, we can reuse them and drop $\{g_{i}\}_{i=1,\ldots,n}\cup\{h\}$. (c) Disturbance-aware Selection: We select the appropriate maintenance for every single temporal feature utilizing the ratio of temporal feature error (i.e., Eq. (14)) between the two maintenance methods. This framework achieves the maintenance of temporal features and yields promising results.
@@ 4.3 Inducement Analyses
TABLE I: FID, sFID and SQNR on LSUN-Bedrooms $256\times 256$ [95] for LDM-4. “Prev” represents BRECQ. Freeze denotes our trial. We conduct the same experiments with more baselines as “Prev” in Sec. B, which further validate our analysis.
@@ 4.4.1 TIB-based Maintenance
Temporal Information Block (TIB). Based on the inducements, it is crucial to meticulously separate the reconstruction and calibration process for each embedding layer and Residual Bottleneck Block to enhance quantized model performance. Considering the unique structure of the UNet, we consolidate all embedding layers and time embed into a unified Temporal Information Block (TIB), which can be denoted as $\{g_{i}\}_{i=1,\ldots,n}\cup\{h\}$ (see Fig. 5 (a)).
Temporal Information-aware Reconstruction. Based on the Temporal Information Block (TIB), we introduce the Temporal Information-aware Reconstruction (TIAR) to address the first inducement. The optimization goal for the block during the reconstruction phase is captured by the following objective:
$\mathop{\arg\,\min}\limits_{\mathcal{S}^{\mathbf{W}}_{g\circ h},\mathcal{Z}^{\mathbf{W}}_{g\circ h}}\sum_{i=1}^{n}\underbrace{\|g_{i}(h(t))-\widehat{g}_{i}(\widehat{h}(t))\|_{F}^{2}}_{{\color[rgb]{0,0,0}\mathcal{L}^{temporal}_{t,i}}},$ (11)
where $\widehat{h}(\cdot)$ and $\widehat{g}_{i}(\cdot)$ represent the quantized counterparts of $h(\cdot)$ and $g_{i}(\cdot)$, respectively. Moreover, $\mathcal{S}^{\mathbf{W}}_{g\circ h}$ and $\mathcal{Z}^{\mathbf{W}}_{g\circ h}$ denote all the quantization scales and zero offsets for every weight of linear operators  [^7] Linear layers and convolutions. in $\{g_{i}\}_{i=1,\ldots,n}\cup\{h\}$, respectively. This reconstruction strategy focuses on adjusting weights to pursue a minimal disturbance for temporal features.
Finite Set Calibration. To address the challenge posed by the wide span of activations within a finite set for the second inducement, we propose Finite Set Calibration (FSC) for activation quantization. This strategy employs $T$ sets of quantization parameters for activation, such as $\{(s_{T},z_{T}),\ldots,(s_{1},z_{1})\}$ for an arbitrary activation $\mathbf{X}$ within embedding layers and time embed. At time-step $t$, the quantization function for the $\mathbf{X}$ can be expressed as:
@@
where $s_{t}$ and $z_{t}$ are the corresponding scale and zero offset. Equipped with FSC, we can obtain $\mathcal{S}^{\mathbf{X}}_{g\circ h}$ and $\mathcal{Z}^{\mathbf{X}}_{g\circ h}$, which refer to all the quantization scales and zero offsets for the activations  [^8] We consider the outputs and inputs of every linear operator like previous works [79, 19]. in $\{g_{i}\}_{i=1,\ldots,n}\cup\{h\}$, respectively. To be noted, the calibration target for these activations is also aligned with the output of the TIB. Besides that, we find that Min-max [62] can achieve satisfactory results with high efficiency (more evidence in Sec. 5.4.2) for range estimation. In Sec. 5.3.1, we perform a detailed analysis to demonstrate that our method incurs a negligible extra cost for inference.
@@ 4.4.2 Cache-based Maintenance
Further considering sample independence and finitude of the temporal feature, the feature associated with each $t$ and $i$ remains constant and can therefore be pre-computed offline. This allows us to directly optimize the quantization parameters for these pre-computed full-precision features and cache the quantized counterparts with the parameters to address the issues. The objective for all $t$ and $i$ can be formulated as follows:
@@
where $s_{t,i}$ and $z_{t,i}$ denote the quantization scale and zero offset for the pre-gained $\mathbf{X}_{t,i}^{temporal}$ (i.e., the output of $g_{i}\circ h$ at $t$). Finally, the obtained quantization parameters can be defined as $\mathcal{S}^{temporal}=\{s_{t,i}\}_{t=1,\ldots,T\&i=1,\ldots,n}$ and $\mathcal{Z}^{temporal}=\{z_{t,i}\}_{t=1,\ldots,T\&i=1,\ldots,n}$. Specifically, in this strategy, we employ LSQ [12] to optimize every objective  [^9] From Eq. (13), it is obvious that we have to optimize $n\times T$ objectives across different indexes of temporal features. separately, which can be done in just a few minutes. During inference, this strategy enables us to get the temporal feature by reloading the cached items. In addition, we find that this approach can help improve latency in cases. More detailed analysis can be found in Sec. 5.3.2.
For clarity, we demonstrate the difference between the two proposed maintenance strategies. In TIB-based Maintenance, we optimize quantization parameters of weights and activations for $\{g_{i}\}_{i=1,\ldots,n}\cup\{h\}$. In contrast, we only adjust the quantization parameters of the final output activations of $\{g_{i}\}_{i=1,\ldots,n}\cup\{h\}$ in Cache-based Maintenance, which implies that $\mathcal{S}^{temporal}\in\mathcal{S}^{\mathbf{X}}_{g\circ h}$ and $\mathcal{Z}^{temporal}\in\mathcal{Z}^{\mathbf{X}}_{g\circ h}$. Thus, it is evident that Eq. (11) and Eq. (13) have the same objective (i.e., minimize $\mathcal{L}_{t,i}^{temoral}$ at corresponding $t$ and $i$), but optimize different sets of parameters in different ways.
@@ 4.4.3 Disturbance-aware Selection
We visualize the comparison of the temporal feature disturbance between the two maintenance approaches in Fig. 6, where the grid value of $(t,i)$ represents the ratio of temporal feature error [^10] Here we employ $\mathcal{L}_{t,i}^{temporal}$ in Eq. (13), instead of that in Eq. (8) as temporal feature error, since over $82\%$ of cosine similarity with our two maintenance approaches are almost exactly equal to 1, which shows the remarkable effect of our methods. between the two approaches as follows:
@@
We believe the discrepancy in temporal feature errors across different $t$ and $i$ between our two methods results from their different sets of optimization parameters and optimization methods. Based on the above observation, we propose Disturbance-aware Selection to better mitigate the temporal feature disturbance. Specifically, when $\tau_{t,i}<1$, we employ TIB-based Maintenance to obtain $\widehat{\mathbf{X}_{t,i}^{temporal}}$. Conversely, we opt for Cache-based Maintenance. The pipeline of our selection is shown in Fig. 5 (c). For efficiency, our selection procedure can be done offline (i.e., before deployment) with only one image generation pass to calculate all the $\tau_{t,i}$. To validate its performance, Fig. 7 demonstrates that our straightforward but effective selection strategy successfully leverages the strengths of both maintenance approaches.
5 Experiments
We organize the experiments as follows: In Sec. 5.1, we first demonstrate implementation details. Then, we exhibit the outstanding performance gained by our framework in Sec. 5.2. Further, we also deploy our quantized models on various hardware to investigate their inference efficiency in Sec. 5.3. Last, we conduct comprehensive ablative studies for the proposed framework in Sec. 5.4. It is worth noting that we also discuss and conduct experiments for video generation in Sec. C. A guideline on when to use which proposed strategies and sub-4-bit quantization results are included in Sec. D and Sec. E, respectively.
@@ 5.2 Performance Comparison
Unconditional generation. In the experiments conducted on the LDM with DDIM sampler [82], we maintain the same experimental settings as presented in  [74], including the number of steps, variance schedule, and classifier-free guidance scale (denoted by eta and cfg in the following, respectively). As shown in Tab. II, the FID performance differences relative to the full precision (FP) model are all within 0.7 for all settings employing DS. Specifically, on the CelebA-HQ $256\times 256$ dataset, DS exhibits an FID reduction of 6.74, an sFID reduction of 6.79, and an SQNR increase of 4.09 in the w4a8 setting compared to the current algorithms. It is noticeable that existing methods, whether in 4-bit or 8-bit, show significant performance degradation when compared to the FP model on face datasets like CelebA-HQ $256\times 256$ and FFHQ $256\times 256$, whereas our TM/CM/DS shows almost no performance degradation. Importantly, DS also achieves significant performance improvement on the LSUN-Bedrooms $256\times 256$ in the w4a8 setting, with FID and sFID reductions of 2.33 and 7.67 compared to PTQD [19], respectively. Regarding LDM-8 on LSUN-Churches $256\times 256$, we attribute the moderate improvement, compared to other datasets. We believe that the use of the LDM-8 model with a downsampling factor of 8 may be more quantization-friendly. Existing methods have already achieved satisfactory results on this dataset. Nonetheless, our method still approaches the performance of the FP model more closely than existing methods.
@@
Besides the experiments on LDM, We have also conducted experiments with DDPM on CIFAR-10 $32\times 32$. As shown in Tab. III, our methods still achieve comprehensive improvements in terms of IS, FID, and SQNR compared to the existing advanced methods. However, due to the lower resolution and simplicity of the images in this dataset, existing methods show minimal performance degradation, so the results we obtain may not be as pronounced.
Class-conditional generation. On the ImageNet $256\times 256$ dataset, we employed a denoising process with 20 steps DDIM sampler [82], setting eta and cfg to 0.0 and 3.0, respectively. In Tab. IV, compared to PTQD, our method achieves a FID reduction of 2.09 on w8a8, and a 5.35 sFID decrease on w4a8. Under all the conditions with DS, we observe an improvement of over 9 in IS. Particularly noteworthy is that, across various quantization settings, our method consistently achieved lower FID and higher SQNR compared to the FP model.
TABLE IV: Quantization results for unconditional image generation with class-conditional image generation with LDM-4 on ImageNet $256\times 256$.
@@
TABLE V: Quantization results for text-guided image generation with Stable Diffusion-v1-4, SD-XL-base-1.0, and SD-XL-turbo on MS-COCO captions.
[…]
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Text-guided generation. In this experiment, we generate high-resolution images of $512\times 512$ pixels using Stable Diffusion-v1-4 with 50 steps PLMS sampler [51], and SD-XL-turbo with only a single step Euler sampler [31]. Moreover, higher resolution images of $1024\times 1024$ are also generated from SD-XL-base-1.0 with 50 steps Euler sampler. cfg is fixed to the default 7.5 for Stable Diffusion and SD-XL as the trade-off between generation quality and diversity. However, we do not use classifier-free guidance [23] to save memory, the same as the original paper [78]. In Tab. V, compared to Q-Diffusion, our approach achieves an FID reduction of 1.20, 5.61, and 5.50 under the w4a8 configuration for Stable Diffusion, SD-XL, and SD-XL-turbo, respectively. Notably, our FID on w8a8 and sFID on w4a32 are even lower than those of the full precision model. The table also shows our significantly lower quantization error than Q-Diffusion. For example, DS obtains an 8.16 SQNR upswing for w4a8 Stable Diffusion. Benefiting from this, the visual effect of our generated images  [^12] More qualitative comparisons can be found in Sec. H. is closer to the FP model with better quality than Q-Diffusion (examples in Fig. 8). All of these clues strongly reveal the value of maintaining temporal features. More experiments for few-step distilled model FLUX.1-Schnell [42] can also be found in Sec. F.